OpenStudy (secret-ninja):

The legs of a right triangle are 9 cm and 40 cm long. What is the length of the hypotenuse? forty one centimeters forty five centimeters forty nine centimeters fifty centimeters

OpenStudy (anonymous):

A(: you need to use pythagorean theorem! ill teach u if u want?

OpenStudy (texaschic101):

a^2 + b^2 = c^2 9^2 + 40^2 = c^2 81 + 1600 = c^2 1681 = c^2 sq rt 1681 = c 41 = c hypotenuse is 41

OpenStudy (secret-ninja):

no thats ok i learned it im just a little rusty lol thanks tho

OpenStudy (anonymous):

np.

OpenStudy (anonymous):

@texaschic101 is A the right answer?

OpenStudy (texaschic101):

yes it is

OpenStudy (anonymous):

thanks ^_^ heres a medal

OpenStudy (anonymous):

medal me pwease

OpenStudy (texaschic101):

no problem ..happt to help

OpenStudy (texaschic101):

oops...typo..happy to help

OpenStudy (anonymous):

lol xD